Circular No. 5702

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. Telephone 617-495-7244/7440/7444 (for emergency use only) TWX 710-320-6842 ASTROGRAM CAM EASYLINK 62794505 MARSDEN@CFA or GREEN@CFA (.SPAN, .BITNET or .HARVARD.EDU)

GRB 930131 J. Ryan, M. Kippen, and M. Varendorff, University of New Hamp-

shire, report on behalf of the Compton Telescope (COMPTEL)/Compton Observatory Team: “The COMPTEL team has imaged a γ-ray burst at energies > 0.72 MeV that took place on Jan. 31.78971 UT. This event was recorded by the BATSE experiment and was noted as being one of the most intense bursts yet observed by BATSE. The burst exhibited two short peaks in intensity, spanning a period of just over 1 s. The COMPTEL analysis (which incorporated a total of 41 events) has resulted in a 1σ source location error box whose corners are located at the following positions (equinox J2000.0): α = 12^{h}13^{m}.0, δ = −7^{o}33′; 12^{h}23^{m}.0, −7^{o}37′; 12^{h}22^{m}.6, −10^{o}26′; 12^{h}12^{m}.6, −10^{o}19′. This error box is roughly 3^{o} on a side and is centered at α = 12^{h}17^{m}.8, δ = −8^{o}59′. We are soliciting collaborators to observe the region with large field-of-view or rastering instruments in an effort to identify a counterpart at other wavelengths. The effort is being coordinated by B. McNamara at New Mexico State University (bmcnamar@nmsu.edu; telephone 505-646-2614) and J. Ryan (jryan@unh.edu; 603-862-3510).”

COMET SHOEMAKER-LEVY (1991a_{1}) No observations of this comet since last August have been reported.

The following ephemeris is from orbital elements on MPC 20602:

1993 TT α_{2000} δ_{2000} Δ r ϵ β m_{1}
Jan. 13 9^{h}58.^{m}1 −70^{o}07.^{′}5 2.715 2.809 85.^{o}1 20.^{o}4 14.2
23 8 57.1 −70 27.7 2.762 2.929 89.9 19.6 14.4
Feb. 2 8 00.70 −68 59.5 2.827 3.049 93.4 18.8 14.6
12 7 18.12 −66 09.4 2.912 3.167 95.7 18.1 14.8
22 6 49.96 −62 33.8 3.018 3.283 96.7 17.4 15.1
Mar. 4 6 32.95 −58 41.5 3.145 3.399 96.2 16.9 15.3
14 6 23.69 −54 51.4 3.289 3.512 94.7 16.4 15.5
24 6 19.70 −51 14.6 3.448 3.625 92.2 16.0 15.8
Apr. 3 6 19.34 −47 57.0 3.619 3.737 89.0 15.5 16.0
13 6 21.49 −45 01.8 3.797 3.847 85.3 15.1 16.2
23 6 25.40 −42 29.7 3.981 3.956 81.3 14.6 16.5
May 3 6 30.56 −40 20.7 4.166 4.064 77.2 14.0 16.7
13 6 36.57 −38 33.8 4.349 4.171 73.2 13.4 16.9

1993 February 2 Daniel W. E. Green